How many ways can two teams of 4 players be chosen from 10 people?
Answer by sudhanshu_kmr(1152)

no. of ways to choose players for first team = 10C4
no. of ways to choose players for second team = 6C4
no. of ways = 210 * 15 = 3150
(if there are two particular team, i.e team A and team B)



if order of team doesn't matter
then no. of ways = 210 * 15 / 2! = 1575